What is the probability that i will choose a face card given that i chose a heart card ?
What is the probability that a couple has three boys and two girls ?
What is the probability that i will choose a face and a heart card ?
What is the probability that i will choose a face or a heart card ?
What is the probability that a couple has two boys and three girls ?

Answer by stanbon(75887)

What is the probability that i will choose a face card given that i chose a heart card ?
Ans: 3/13 because there are 3 hearts that are face cards (K,Q,and J)
------------------------------------------------------------------------
What is the probability that a couple has three boys and two girls ?
Ans: 5C3*(1/2)^3*(1/2)^2 = 10/32
-----------------------------------------
What is the probability that i will choose a face and a heart card ?
Do you mean a single card that is a face card and a heart?
If so, the answer is 3/52 = 1/14
--------------------------------------------
What is the probability that i will choose a face or a heart card ?
P(face OR heart) = P(face) + P(heart) - P(face AND heart)
= (12/52) + 13/52 - 3/52 = 22/52
------------------------------------------
What is the probability that a couple has two boys and three girls ?
Ans: 5C2*(1/2)^2*(1/2)^3 = 10/32
============================================
Cheers,
Stan H.
==================